OpenStudy (anonymous):

7. Write an equation in slope intercept form of the line that passes through (4, 1) and (5, -1).(Points : 1) y = -4x + 1 y = -1/2 x + 9 y = 4x + 1 y = -2x + 9 8. Write an equation of the line that passes through (8, 3) and (4, 3).(Points : 1) y = 3 y = 4x + 3 y = 4 y = 3x

OpenStudy (blurbendy):

(-1 - 1) = -2 (5 -4 ) = 1 y = -2x +9 8) (3 -3) = 0 (4 - 8) = -4 y = 3